# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

### v4

@@ 2.6 Local adapted gauge and semicontinuity argument
Theorem 2.18 (Semicontinuity of off-diagonal holonomy dimension).
The function $\varepsilon\mapsto\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower semicontinuous on $[0,1]$. In particular,
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq\liminf_{\varepsilon\to 0^{+}}\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}}).$
Proof.
The curvature tensor $R^{C_{\varepsilon}}$ depends smoothly on $\varepsilon$ and on the point $p\in M$. At each $p$, the off-diagonal holonomy algebra is generated by the image of the linear map $\Phi_{\varepsilon}:\bigwedge^{2}T_{p}M\to\mathfrak{so}(T_{p}M)$ given by $\Phi_{\varepsilon}(X\wedge Y)=R^{C_{\varepsilon}}(X,Y)$. The dimension of the Lie algebra generated by $\operatorname{im}\Phi_{\varepsilon}$ is lower semicontinuous in $\varepsilon$ because the rank of a continuous family of linear maps is lower semicontinuous. ∎
2.7 Summary of the geometric setup for the PT lower bound
We have established the following geometric framework for applying the PT lower bound:
@@ 2.7 Summary of the geometric setup for the PT lower bound
    The extended parameter space is the product manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, equipped with a family of metrics $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ interpolating between a Riemannian product metric ($\varepsilon=0$) and the physical Kaluza–Klein metric ($\varepsilon=1$).
@@
    For each $\varepsilon$, we have a metric connection $\nabla^{C_{\varepsilon}}$ with totally skew-symmetric torsion $T_{\varepsilon}$. The cohomology class $T_{\varepsilon}]$ is constant in $\varepsilon$ and has a specific Künneth decomposition.
@@
    The PT theorem applies directly to the product case $\varepsilon=0$, giving a lower bound on the off-diagonal holonomy dimension in terms of the mixed rank and the obstruction kernel $\mathcal{K}_{0}$.
@@
    For $\varepsilon>0$, the obstruction kernel $\mathcal{K}_{\varepsilon}$ vanishes, but the dimension of the space of vertical parallel 1-forms drops from 2 to 1. This discontinuity is carefully handled by the semicontinuity argument.
  5. 5.
    The semicontinuity of the off-diagonal holonomy dimension allows us to transfer the lower bound from $\varepsilon=0$ to $\varepsilon=1$, provided we can control the behavior at $\varepsilon=0$.
In the next section, we compute the topological invariants (mixed rank and obstruction kernel) explicitly for our geometry and apply the PT theorem to the product metric $g_{0}$.
@@ 3.1 De Rham cohomology and Künneth decomposition
$H^{3}(M;\mathbb{R})\cong$ $\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})\bigr)$
$\oplus\bigl(H^{1}(T^{2}_{\mathrm{BZ}})\otimes H^{2}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})\bigr).$
@@ 3.2 Harmonic representative for the product metric
$\omega_{h}^{0}=\operatorname{vol}_{\mathrm{BZ}}\wedge\bigl(a_{+}d\phi_{+}+a_{-}d\phi_{-}\bigr),$
@@
Since $g_{0}$ is a product metric, the Hodge Laplacian decouples. The form $\eta:=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}d\phi_{+}+a_{-}d\phi_{-})$ is closed because $\operatorname{vol}_{\mathrm{BZ}}$ is closed and $d\phi_{\pm}$ are closed. It is coclosed iff
$\delta\eta=-*d*\eta=0.$
In the product metric, $\eta=a_{+}*_{\mathrm{BZ}}\operatorname{vol}_{\mathrm{BZ}}\wedge*_{\text{fibre}}d\phi_{+}+a_{-}*_{\mathrm{BZ}}\operatorname{vol}_{\mathrm{BZ}}\wedge*_{\text{fibre}}d\phi_{-}$. Since $_{\mathrm{BZ}}\operatorname{vol}_{\mathrm{BZ}}=1$ and $_{\text{fibre}}d\phi_{\pm}$ are constant, we obtain the condition $a_{+}c^{(+)}+a_{-}c^{(-)}=0$, i.e., $a_{-}=\lambda a_{+}$. Uniqueness follows from the Hodge theorem. ∎
@@ 3.3 Mixed tensor rank
$\operatorname{rank}_{\mathbb{R}}([\eta]):=\min\bigl\{r_{2,1}+r_{1,2}\bigr\},$
@@
$\operatorname{rank}_{\mathbb{R}}([\omega])=1.$
@@
From the explicit form of the harmonic representative $\omega_{h}^{0}$, we have
$\omega]=[\operatorname{vol}_{\mathrm{BZ}}]\otimes\bigl(a_{+}[d\phi_{+}]+a_{-}[d\phi_{-}]\bigr).$
Since $a_{+}[d\phi_{+}]+a_{-}[d\phi_{-}]$ is a single element in $H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, this shows that the mixed rank is at most 1. It cannot be 0 because $\omega]\neq 0$. ∎
@@ 3.4 Parallel forms and obstruction kernel for the product metric
Let $\mathcal{V}_{2,1}$ be the subspace of $H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ spanned by the harmonic representatives of the $(2,1)$-component of $\omega]$ with respect to $g_{0}$. Define
$\mathcal{K}_{0}:=\bigl(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\bigr)\oplus\bigl(\mathcal{V}_{1,2}\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\bigr).$
@@
Thus $\mathcal{V}_{2,1}=\operatorname{span}\{\xi_{0}\}$. Since $\operatorname{vol}_{\mathrm{BZ}}\in\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ and $a_{+}d\phi_{+}+a_{-}d\phi_{-}\in\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ (it is a constant linear combination), we have $\xi_{0}\in\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$. Hence
$\mathcal{V}_{2,1}\subset\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}),$
so the intersection is all of $\mathcal{V}_{2,1}$, and $\dim\mathcal{K}_{0}=\dim\mathcal{V}_{2,1}=1$. The $(1,2)$-component is zero, so it contributes nothing. ∎
@@ 3.5 Reduced rank for the product metric
$r^{\sharp}_{0}:=\operatorname{rank}_{\mathbb{R}}([\omega])-\dim\mathcal{K}_{0}.$
Corollary 3.12 (Value of $r^{\sharp}_{0}$).
@@
By Theorem 3.6, $\operatorname{rank}_{\mathbb{R}}([\omega])=1$. By Theorem 3.10, $\dim\mathcal{K}_{0}=1$. Hence $r^{\sharp}_{0}=1-1=0$. ∎
@@
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r^{\sharp}_{0}=0,$
@@ 3.6 Behavior under deformation
For $\varepsilon>0$, the space $\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ consists of vertical 1-forms $\eta=c_{+}d\phi_{+}+c_{-}d\phi_{-}$ satisfying $c_{+}+\lambda c_{-}=0$ (Theorem 2.12). The harmonic representative of the $(2,1)$-component with respect to $g_{\varepsilon}$ is not a simple tensor product of forms on the base and fibre, but its projection onto $\mathcal{V}_{2,1}$ in the Künneth decomposition is still $\xi_{0}=\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})$. However, now $a_{+}d\phi_{+}+a_{-}d\phi_{-}$ does not belong to $\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ because it does not satisfy the condition $a_{+}+\lambda a_{-}=0$ (instead, it satisfies $a_{-}=\lambda a_{+}$). Indeed,
$a_{+}+\lambda a_{-}=a_{+}+\lambda(\lambda a_{+})=a_{+}(1+\lambda^{2})\neq 0.$
Thus $\xi_{0}\notin\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$. Hence the intersection is trivial. The $(1,2)$-component remains zero. ∎
@@
For $\varepsilon>0$, the reduced rank $r^{\sharp}_{\varepsilon}:=\operatorname{rank}_{\mathbb{R}}([\omega])-\dim\mathcal{K}_{\varepsilon}$ satisfies
$r^{\sharp}_{\varepsilon}=1.$
Thus, while the PT theorem applied directly to $g_{0}$ gives a trivial bound, the geometry for $\varepsilon>0$ has a non-trivial reduced rank. The key step is to use semicontinuity to transfer the bound from $\varepsilon>0$ to $\varepsilon=1$.
@@ 3.7 Summary
    The torsion class $\omega]\in H^{3}(M;\mathbb{R})$ has mixed tensor rank 1 (Theorem 3.6), a topological invariant.
@@
    For the product metric $g_{0}$, the obstruction kernel $\mathcal{K}_{0}$ has dimension 1, giving reduced rank $r^{\sharp}_{0}=0$ (Corollary 3.12).
@@
    For the deformed metrics $g_{\varepsilon}$ with $\varepsilon>0$, the obstruction kernel vanishes, giving reduced rank $r^{\sharp}_{\varepsilon}=1$ (Corollary 3.15).
@@
    $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r^{\sharp}_{0}=0.$
    This bound is trivial, but the semicontinuity argument developed below will allow us to obtain a non-trivial bound for $\varepsilon=1$.
In the next section, we adapt the proof of the PT theorem to our deformation family and use semicontinuity to establish the lower bound for the physical connection $\nabla^{C}$.
4 Applying the PT Theorem via Deformation and Semicontinuity
@@
    Apply the original PT theorem to the product end $(\varepsilon=0)$ of the deformation.
@@
    Show by a direct analysis that the lower bound improves for $\varepsilon>0$.
@@
    Use lower semicontinuity to transfer the non‑trivial bound to the physical endpoint $\varepsilon=1$.
@@ 4.1 Applying the PT theorem to the product metric $g_{0}$
From Corollary 3.12 we have $r^{\sharp}_{0}=\operatorname{rank}_{\mathbb{R}}([\omega])-\dim\mathcal{K}_{0}=1-1=0$. ∎
Remark 4.2.
@@
The bound is trivial for the product metric, as expected. In the product geometry the connection $\nabla^{C_{0}}$ may well have block‑diagonal curvature; the topological obstruction encoded by the mixed class is completely absorbed by the parallel forms of the product metric (see Remark 2.15).
4.2 Direct analysis for the deformed metrics $\varepsilon>0$
For $\varepsilon>0$ the metric $g_{\varepsilon}$ is no longer a product, and the spaces of parallel forms change discontinuously at $\varepsilon=0$ (Theorem 2.12). Consequently the PT theorem does not apply directly. Nevertheless, a pointwise computation in adapted gauge reveals that the off‑diagonal curvature acquires a non‑trivial generator.
Definition 4.3 (Pointwise obstruction map for $\varepsilon>0$).
Fix $\varepsilon>0$ and a point $p\in M$. Choose a gauge adapted at $p$ (Lemma 2.17). Let
$\xi_{0}=\operatorname{vol}_{\mathrm{BZ}}\otimes(d\phi_{+}+\lambda d\phi_{-}),\qquad\lambda=c^{(-)}/c^{(+)}.$
Define $\Psi_{p}^{(\varepsilon)}:\operatorname{span}\{\xi_{0}\}\to\mathfrak{so}(T_{p}M)$ by
$\Psi_{p}^{(\varepsilon)}(t\xi_{0}):=\frac{1}{2}\bigl(\nabla^{LC(\varepsilon)}T_{\varepsilon}\bigr)_{\mathrm{off}}(X,Y,Z),\qquad t\in\mathbb{R},$
where $X,Y\in\mathcal{H}_{p}^{(\varepsilon)}$, $Z\in\mathcal{V}_{p}$ satisfy $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$ and $(d\phi_{+}+\lambda d\phi_{-})(Z)=1$, and $(\cdot)_{\mathrm{off}}$ extracts the off‑diagonal component (mapping $\mathcal{H}_{p}^{(\varepsilon)}$ to $\mathcal{V}_{p}$ or vice‑versa).
Theorem 4.4 (Injectivity for $\varepsilon>0$).
Under Assumption 2.3, for every $\varepsilon>0$ and every $p\in M$, the map $\Psi_{p}^{(\varepsilon)}$ is injective and satisfies
$\Psi_{p}^{(\varepsilon)}(\xi_{0})=\frac{1}{2}\,a_{+}c^{(+)}(1+\lambda^{2})\,J^{(\varepsilon)}\neq 0,$
where $a_{+}$ is the constant appearing in the harmonic representative of $\omega]$ (Theorem 3.3) and $J^{(\varepsilon)}:T_{p}M\to T_{p}M$ is a fixed off‑diagonal operator exchanging horizontal and vertical directions.

### v5

@@ 2.6 Local adapted gauge
2.7 Summary of the geometric framework
We have established the following geometric framework:
@@ 2.7 Summary of the geometric framework
    The extended parameter space is the smooth product manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, equipped with a one-parameter family of metrics $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ interpolating between a Riemannian product metric ($\varepsilon=0$) and the physical Kaluza–Klein metric ($\varepsilon=1$).
@@
    For each $\varepsilon$, we have a metric connection $\nabla^{C_{\varepsilon}}$ with totally skew-symmetric torsion $T_{\varepsilon}$. The cohomology class $T_{\varepsilon}]$ is constant in $\varepsilon$ and admits a Künneth decomposition into mixed components (Lemma 2.9).
@@
    The space of vertical parallel 1-forms undergoes a discontinuous jump at $\varepsilon=0$: $\dim\mathcal{P}_{1}^{0}=2$ while $\dim\mathcal{P}_{1}^{\varepsilon}=1$ for all $\varepsilon>0$ (Theorem 2.13 and Lemma 2.14).
@@
    Consequently, the obstruction kernel $\mathcal{K}_{\varepsilon}$ vanishes for all $\varepsilon>0$ (Lemma 2.16), while $\mathcal{K}_{0}$ is non-trivial. This discontinuity is the key to obtaining a non-trivial lower bound for the physical connection.
In the next section, we compute the topological invariants (mixed cohomology class and its tensor rank) explicitly and determine the reduced rank $r^{\sharp}_{\varepsilon}=\operatorname{rank}_{\mathbb{R}}([T_{\varepsilon}]_{\mathrm{mixed}})-\dim\mathcal{K}_{\varepsilon}$ for both $\varepsilon=0$ and $\varepsilon>0$.
@@ 3.1 De Rham cohomology and Künneth decomposition
$H^{3}(M;\mathbb{R})\cong\left(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})\right)\oplus\left(H^{1}(T^{2}_{\mathrm{BZ}})\otimes H^{2}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})\right).$
@@ 3.2 Harmonic representative for the product metric
$\omega_{h}^{0}=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}d\phi_{+}+a_{-}d\phi_{-}),$
@@
Since $g_{0}$ is a product metric, the Hodge Laplacian decouples. The form $\eta:=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}d\phi_{+}+a_{-}d\phi_{-})$ is closed because $\operatorname{vol}_{\mathrm{BZ}}$ is closed and $d\phi_{\pm}$ are closed.
It is coclosed if and only if $\delta\eta=0$. In the product metric, using $\eta=a_{+}(*_{\mathrm{BZ}}\operatorname{vol}_{\mathrm{BZ}})\wedge(*_{\text{fibre}}d\phi_{+})+a_{-}(*_{\mathrm{BZ}}\operatorname{vol}_{\mathrm{BZ}})\wedge(*_{\text{fibre}}d\phi_{-})$, where $_{\mathrm{BZ}}\operatorname{vol}_{\mathrm{BZ}}=1$ and $_{\text{fibre}}d\phi_{\pm}=\pm d\phi_{\mp}$, the coclosed condition gives:
$a_{+}c^{(+)}+a_{-}c^{(-)}=0,$
which is equivalent to $a_{-}=\lambda a_{+}$. Uniqueness follows from the Hodge theorem. ∎
@@ 3.3 Mixed tensor rank
$\operatorname{rank}_{\mathbb{R}}([\eta]_{\mathrm{mixed}}):=\min\{r_{2,1}+r_{1,2}\},$
@@
$\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})=1.$
@@
From the explicit form of the harmonic representative $\omega_{h}^{0}$ (Theorem 3.3), we have
$\omega]=[\operatorname{vol}_{\mathrm{BZ}}]\otimes\left(a_{+}[d\phi_{+}]+a_{-}[d\phi_{-}]\right).$
Since $a_{+}[d\phi_{+}]+a_{-}[d\phi_{-}]$ is a single element in $H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, this shows that the mixed rank is at most 1. It cannot be 0 because $\omega]\neq 0$ (under Assumption 2.3, $c^{(\pm)}\neq 0$). ∎
@@ 3.4 Parallel forms and obstruction kernel for the product metric
Let $\mathcal{V}_{2,1}$ be the subspace of $H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ spanned by the harmonic representatives of the $(2,1)$-component of $\omega]$ with respect to $g_{0}$ (as given in Definition 3.5). Define
@@
Thus $\mathcal{V}_{2,1}=\operatorname{span}\{\xi_{0}\}$. Since $\operatorname{vol}_{\mathrm{BZ}}\in\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ and $a_{+}d\phi_{+}+a_{-}d\phi_{-}\in\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ (it is a constant linear combination), we have
$\xi_{0}\in\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}).$
Hence
$\mathcal{V}_{2,1}\subseteq\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}),$
so the intersection is all of $\mathcal{V}_{2,1}$, and $\dim\mathcal{K}_{0}=\dim\mathcal{V}_{2,1}=1$. The $(1,2)$-component is zero (since $b_{1}(T^{2}_{\mathrm{BZ}})\cdot b_{2}(S^{1}\times S^{1})=0$), so it contributes nothing. ∎
@@ 3.5 Reduced rank for the product metric
$r_{0}^{\sharp}:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}_{0}.$
Corollary 3.12 (Value of $r_{0}^{\sharp}$).
@@
By Theorem 3.6, $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})=1$. By Theorem 3.10, $\dim\mathcal{K}_{0}=1$. Hence $r_{0}^{\sharp}=1-1=0$. ∎
@@
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r_{0}^{\sharp}=0,$
@@ 3.6 Behavior under deformation
The $(2,1)$-component of the torsion class is represented by the simple tensor $\xi_{0}=[\operatorname{vol}_{\mathrm{BZ}}]\otimes(a_{+}[d\phi_{+}]+a_{-}[d\phi_{-}])$ (Theorem 3.3). The kernel $\mathcal{K}_{\varepsilon}$ is defined by the intersection of the span of its harmonic representative with the space $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$.
As established in Theorem 2.13, for $\varepsilon>0$, any parallel vertical 1-form $\eta\in\mathcal{P}_{1}^{\varepsilon}$ must satisfy the algebraic constraint $c_{+}+\lambda c_{-}=0$, where $\lambda=c^{(-)}/c^{(+)}$. However, the harmonic representative of $\omega]$ has coefficients satisfying $a_{-}=\lambda a_{+}$. Substituting these into the parallel condition yields:
$a_{+}+\lambda a_{-}=a_{+}+\lambda(\lambda a_{+})=a_{+}(1+\lambda^{2}).$
Since $c^{(\pm)}\neq 0$ (Assumption 2.3), it follows that $1+\lambda^{2}>0$ for any real $\lambda$. Thus, the parallel condition $a_{+}(1+\lambda^{2})=0$ is satisfied if and only if $a_{+}=0$, which implies $\omega]=0$. For any non-trivial torsion class, the intersection is therefore $\{0\}$, proving that $\mathcal{K}_{\varepsilon}=0$. ∎
@@
For $\varepsilon>0$, the reduced rank $r_{\varepsilon}^{\sharp}:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}_{\varepsilon}$ satisfies
$r_{\varepsilon}^{\sharp}=1.$
Proof.
By Theorem 3.6, $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})=1$. By Theorem 3.14, $\dim\mathcal{K}_{\varepsilon}=0$ for $\varepsilon>0$. Hence $r_{\varepsilon}^{\sharp}=1-0=1$. ∎
@@ 3.7 Summary
    The torsion class $\omega]\in H^{3}(M;\mathbb{R})$ has mixed tensor rank $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})=1$ (Theorem 3.6), a topological invariant.
@@
    For the product metric $g_{0}$, the obstruction kernel $\mathcal{K}_{0}$ has dimension 1, giving reduced rank $r_{0}^{\sharp}=0$ (Corollary 3.12).
@@
    For the deformed metrics $g_{\varepsilon}$ with $\varepsilon>0$, the obstruction kernel vanishes, giving reduced rank $r_{\varepsilon}^{\sharp}=1$ (Corollary 3.15).
@@
    $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r_{0}^{\sharp}=0.$
    This bound is trivial, but for $\varepsilon>0$ the bound becomes non-trivial, as will be established in Section 4 through direct curvature analysis.
In the next section, we apply the deformation strategy and establish the non-trivial lower bound for the physical connection $\nabla^{C}=\nabla^{C_{1}}$.
4 Applying the PT Theorem via Deformation and Direct Analysis
@@
    Apply the original PT theorem to the product metric $g_{0}$ at $\varepsilon=0$.
@@
    Establish the lower bound for all $\varepsilon>0$ through direct curvature analysis in adapted gauge, working with the torsion form explicitly.
@@
    Apply the result to the physical connection at $\varepsilon=1$.
The key observation is that while the PT bound is trivial at the product limit ($\varepsilon=0$), it becomes non-trivial for all $\varepsilon>0$ due to the vanishing of the obstruction kernel (Lemma 2.16). This allows us to establish the desired inequality for the physical system through explicit parametric calculation in the constants $c^{(\pm)}$ characterizing the SOC BEC model.
@@ 4.1 Applying the PT theorem to the product metric $g_{0}$
From Corollary 3.12 (Section 3) we have
$r^{\sharp}_{0}=\min(\operatorname{rank}_{\mathbb{R}}([\omega])-\dim\mathcal{K}_{0},1)=\min(1-1,1)=0.$
Remark 4.2 (Triviality at the product limit).
@@
The bound is trivial for the product metric, as expected. In the product geometry, the connection $\nabla^{C_{0}}$ may have block-diagonal curvature because the topological obstruction encoded by the mixed class is completely absorbed by the parallel forms: the harmonic representative of $\omega]$ belongs to $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, yielding $\dim\mathcal{K}_{0}=1$ and hence $r^{\sharp}_{0}=0$. This absorption is the geometric reason for the vanishing of the lower bound at $\varepsilon=0$.
4.2 Direct curvature analysis for $\varepsilon>0$
For $\varepsilon>0$, the metric $g_{\varepsilon}$ is no longer a Riemannian product, and the PT theorem does not apply directly. However, we can establish the lower bound through explicit pointwise computation of the off-diagonal curvature components for the SOC BEC model.
The key geometric fact is that for $\varepsilon>0$, the space of vertical parallel 1-forms drops from dimension 2 to dimension 1 (Theorem 2.13), causing the obstruction kernel to vanish: $\mathcal{K}_{\varepsilon}=0$ (Lemma 2.16). This yields a non-trivial reduced rank $r^{\sharp}_{\varepsilon}=1$ for all $\varepsilon>0$.
Definition 4.3 (Off-diagonal curvature test for $\varepsilon>0$).
Fix $\varepsilon>0$ and a point $p\in M$. Let $\mathcal{H}_{p}$ and $\mathcal{V}_{p}$ denote the horizontal and vertical subspaces of $T_{p}M$, respectively, and let
$\pi_{\mathrm{off}}:\mathfrak{so}(T_{p}M)\;\longrightarrow\;\mathrm{Hom}(\mathcal{H}_{p},\mathcal{V}_{p})\oplus\mathrm{Hom}(\mathcal{V}_{p},\mathcal{H}_{p})$
be the projection onto the off-diagonal component (endomorphisms that map $\mathcal{H}_{p}\to\mathcal{V}_{p}$ and $\mathcal{V}_{p}\to\mathcal{H}_{p}$).
Define the off-diagonal curvature test at $p$ as the collection of endomorphisms
$\mathcal{R}^{\mathrm{off}}_{p}(\varepsilon)\;:=\;\bigl\{\,\pi_{\mathrm{off}}\!\bigl(R^{C_{\varepsilon}}(X,Z)\bigr)\;\big|\;X\in\mathcal{H}_{p},\;Z\in\mathcal{V}_{p}\,\bigr\}\;\subset\;\mathrm{Hom}(\mathcal{H}_{p},\mathcal{V}_{p})\oplus\mathrm{Hom}(\mathcal{V}_{p},\mathcal{H}_{p}).$
The off-diagonal curvature is non-trivial at $p$ if $\mathcal{R}^{\mathrm{off}}_{p}(\varepsilon)\neq\{0\}$, i.e. if at least one mixed-input curvature operator has a nonzero off-diagonal projection.
Remark 4.4 (Intrinsic nature of the off-diagonal curvature test).
Although explicit computations are performed in an adapted gauge (Definition 2.18), the set $\mathcal{R}^{\mathrm{off}}_{p}(\varepsilon)$ is intrinsically defined because:
  1. 1.
    The curvature tensor $R^{C_{\varepsilon}}(X,Z)$ is a $(1,3)$-tensor, independent of any choice of coordinates or gauge.
  2. 2.
    The splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$ is determined by the Kaluza–Klein structure: $\mathcal{V}_{p}=\ker d\pi|_{p}$ and $\mathcal{H}_{p}=\mathcal{V}_{p}^{\perp_{g_{\varepsilon}}}$.
  3. 3.
    The projection $\pi_{\mathrm{off}}$ depends only on the splitting, not on the choice of basis within each factor.
By the Ambrose–Singer theorem, if $\mathcal{R}^{\mathrm{off}}_{p}(\varepsilon)\neq\{0\}$, then $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}})\neq\{0\}$ and hence $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}})\geq 1$.
Theorem 4.5 (Off-diagonal holonomy bound for SOC BECs).
Consider a two-component SOC BEC with extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ equipped with the Kaluza–Klein metric $g_{\varepsilon}$ for $\varepsilon>0$.
Under Assumption 2.3 (constant Berry curvatures $F^{(\pm)}=c^{(\pm)}\operatorname{vol}_{\mathrm{BZ}}$ with $c^{(\pm)}\neq 0$), the off-diagonal holonomy dimension satisfies:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1\quad\text{at every point }p\in M.$ (4.1)